OpenStudy (firejay5):

Solve each system of equations by using substitution: Show work and steps of how you did it. #25 I am trying to find the x value 25. x + 4y = -8 3x + 2y = 6 26. 10x - 9y = 15 5x - 4y = 10 27. 3x - 7y = -3 2x + 6y = -34

OpenStudy (anonymous):

for each of the equation times it by a -1 then subtract the 2 equations from each other.

OpenStudy (anonymous):

I will do one of them, then you do the rest

OpenStudy (firejay5):

I mean I have y = 3, I just need to put it into one of the equations; which I need help with

OpenStudy (anonymous):

ok, but all of the 3 equations that you have will have seperate x and y values, so we need to know from which equation did you derive that y=3

OpenStudy (firejay5):

#25

OpenStudy (anonymous):

ok, so then plug it into equation 25 to get x. use this to solve for x (x + 4y = -8)

OpenStudy (firejay5):

ok

OpenStudy (anonymous):

once you get the values x and y make sure that when you input both of them into equation 25 it makes sense. plug them into either equation of #25

OpenStudy (firejay5):

I mean plug it in

OpenStudy (anonymous):

wait, let me see first if your y is correct

OpenStudy (firejay5):

3(-4y -8) + 2y = 6 -12y - 24 + 2y = 6 -10y -24 = 6 add 24 to both sides -10y = 30 divide by -10y to both sides y = 3

OpenStudy (anonymous):

no, y=-3, your steps are correct but not the last one.

OpenStudy (firejay5):

oh sorry I meant to put -3

OpenStudy (anonymous):

its all good, yeah thats where maths can play up, lol

OpenStudy (firejay5):

okay now can you show me y = -3 being plugged in

OpenStudy (anonymous):

ok sure, x+4(-3)=-8 x-12=-8 x=-8+12 x=4

OpenStudy (firejay5):

thanks @heedcom I am a visual learner btw. If I don't see it, I won't learn anything.

OpenStudy (anonymous):

now check that both the y and the x values are correct by plugging them into the other equation of #25 you should get 6

OpenStudy (firejay5):

yep it works

OpenStudy (anonymous):

when it comes to maths we are all visual learners so the same way, you should be able to do the others, I will remain here, let me know if your stuck

OpenStudy (firejay5):

so will you be on my question or others

OpenStudy (anonymous):

oh dont worry, i will remain on this question

OpenStudy (firejay5):

May I ask you this, is #26 a solution

OpenStudy (anonymous):

what do you mean is #26 a solution

OpenStudy (firejay5):

can't we a no solution sometimes for substitution

OpenStudy (anonymous):

correct, but I would doubt that they would do that, no solution is only when it x or y=0 let me try, wait

OpenStudy (firejay5):

@heedcom quick question

OpenStudy (anonymous):

found them, both the x and y values.

OpenStudy (anonymous):

yes, whats the quick question

OpenStudy (firejay5):

The math book says #17 is a no solution: 17. 6x - 3y = 12 2x = 8 - y

OpenStudy (anonymous):

ok, wait, let me try

OpenStudy (firejay5):

thanks

OpenStudy (anonymous):

well it seems like it does, I have just found the y value, but wait I will find x

OpenStudy (anonymous):

it does have a solution!!!!!!!!!

OpenStudy (firejay5):

I mean I have asked other people is it's a solution or not a solution, and I want to make sure

OpenStudy (anonymous):

it is a solution, but you said it was the book that said that, now your saying its other people!!

OpenStudy (firejay5):

I have asked people if it was a solution or not a solution

OpenStudy (anonymous):

its all good, so are you good with solving them now

OpenStudy (anonymous):

i can give you the steps for the second one, but thats all I can, as I have studies to get through

OpenStudy (firejay5):

nah don't worry about it

OpenStudy (anonymous):

ok, I think you have got it